# Long Context Transfer from Language to Vision

### v1

@@ Abstract
Video sequences offer valuable temporal information, but existing large multimodal models (LMMs) fall short in understanding extremely long videos. Many works address this by reducing the number of visual tokens using visual resamplers. Alternatively, in this paper, we approach this problem from the perspective of the language model. By simply extrapolating the context length of the language backbone, we enable LMMs to comprehend orders of magnitude more visual tokens without any video training. We call this phenomenon long context transfer and carefully ablate its properties. To effectively measure LMMs’ ability to generalize to long contexts in the vision modality, we develop V-NIAH (Visual Needle-In-A-Haystack), a purely synthetic long vision benchmark inspired by the language model’ s NIAH test. Our proposed Long Video Assistant (LongVA) can process 2000 frames or over 200K visual tokens without additional complexities. With its extended context length, LongVA achieves state-of-the-art performance on Video-MME among 7B-scale models by densely sampling more input frames. Our work is open-sourced at https://github.com/EvolvingLMMs-Lab/LongVA.
@@ 1 Introduction
Given the circumstance, in this paper, instead of reducing the visual tokens, we identify the more critical issue limiting the visual context length in existing LMMs: the context length of the language model backbone. Given a language model, we first extend its context length by training on longer text data. We then use this context-extended LM as the backbone to perform modality alignment and visual instruction tuning without any long video text pairs. By training this way, the context length of the language model is directly transferred to that of the LMMs. We further proposed UniRes, a unified encoding scheme that represents videos as extended images, enhancing the capability fusion between images and videos. To facilitate benchmarking and accurately assess the context length in the visual domain, we created V-NIAH, a synthetic visual benchmark based on the Needle-in-a-haystack test [23] used in language models. Our model, Long Video Assistant (LongVA), is capable of accurately retrieving visual information from 2000 frames or more than 200K visual tokens. Experiments show that additional frames during inference lead to improved performance on long video question-answering benchmarks, and LongVA achieves state-of-the-art performance among 7B models on the Video-MME dataset [21]. In summary, our paper makes the following contributions:
@@
(3) Long Video Assistant (LongVA): With long context transfer and UniRes, we developed LongVA that can perceive more than 200K visual tokens, achieving SoTA performance on the Video-MME dataset.
@@ 2 Related Work
Recent years have witnessed significant progress in Video Question-Answering[3]. To accurately measure the progress of the video LMMs’ performance, researchers have developed various benchmarks encompassing a broad spectrum of tasks. These range from fundamental visual perception tasks such as activity recognition[80], concept detection [76], and counting [27], to more complex visual reasoning tasks including compositional [24], causal [72, 79, 77], and situated reasoning [70]. However, most of those benchmarks focus on short videos, lacking data and metrics to test LMMs’ capability over a long context. Inspired by the NIAH test [23] in the language model community, we proposed V-NIAH to benchmark LMMs’ ability over long visual inputs with minimum overhead of data collection and human annotation. Several concurrent works also developed multimodal versions of the Needle-in-a-haystack test [69, 86, 86]. However, they only measure on several hundreds of frames and lack a strong baseline to properly analyze the properties of visual context length.
@@ 3.2 Aligning Long Language Model Using Short Vision Data
Inspired by the AnyRes encoding scheme in LLaVA-NeXT [46, 34], we designed UniRes that provides a unified encoding scheme for both images and videos, as shown in Figure 2. Unlike AnyRes which retains a small base image and flattens ViT patches across the grids, UniRes removes the base image, flattens patches within each grid, and 2x2 pool the visual features by default (Appendix C). This approach allows us to maintain consistent representation when extending image data into videos where multiple frames are viewed as multiple grids in a row.
@@ 5 Experiments
We primarily assess the long visual capability of LongVA on two benchmarks: V-NIAH (Section 5.1 and Video-MME [21] (Section 5.2). V-NIAH provides quick signals about the visual context length of LongVA. Being synthetic, constructing such benchmarks is generally low-cost and can extend to arbitrary context lengths without extra human annotation or data collection. However, it only tests the model’s ability to retrieve information and does not cover other abilities necessary for a real-world long video assistant. Therefore, we also include LongVA’s performance on Video-MME, a comprehensive evaluation suite for video LMMs that includes diverse data types and qualitative annotations. Video-MME is an ideal benchmark for assessing LMMs’ ability to handle long videos in real-world scenarios, given its average video duration of 1017 seconds and the inclusion of short, medium, and long subsets.
@@ Appendix A Needle In A Haystack Test
Appendix B Visual Needle In A Haystack Test

### v2

@@ Abstract
Video sequences offer valuable temporal information, but existing large multimodal models (LMMs) fall short in understanding extremely long videos. Many works address this by reducing the number of visual tokens using visual resamplers. Alternatively, in this paper, we approach this problem from the perspective of the language model. By simply extrapolating the context length of the language backbone, we enable LMMs to comprehend orders of magnitude more visual tokens without any video training. We call this phenomenon long context transfer and carefully ablate its properties. To effectively measure LMMs’ ability to generalize to long contexts in the vision modality, we develop V-NIAH (Visual Needle-In-A-Haystack), a purely synthetic long vision benchmark inspired by the language model’ s NIAH test. Our proposed Long Video Assistant (LongVA) can process 2000 frames or over 200K visual tokens without additional complexities. With its extended context length, LongVA achieves state-of-the-art performance on Video-MME and MLVU among 7B-scale models by densely sampling more input frames. Our work is open-sourced at https://github.com/EvolvingLMMs-Lab/LongVA.
@@ 1 Introduction
Given the circumstance, in this paper, instead of reducing the visual tokens, we identify the more critical issue limiting the visual context length in existing LMMs: the context length of the language model backbone. Given a language model, we first extend its context length by training on longer text data. We then use this context-extended LM as the backbone to perform modality alignment and visual instruction tuning without any long video text pairs. By training this way, the context length of the language model is directly transferred to that of the LMMs. We further proposed UniRes, a unified encoding scheme that represents videos as extended images, enhancing the capability fusion between images and videos. To facilitate benchmarking and accurately assess the context length in the visual domain, we created V-NIAH, a synthetic visual benchmark based on the Needle-in-a-haystack test [23] used in language models. Our model, Long Video Assistant (LongVA), is capable of accurately retrieving visual information from 2000 frames or more than 200K visual tokens. Experiments show that additional frames during inference lead to improved performance on long video question-answering benchmarks, and LongVA achieves state-of-the-art performance among 7B models on the Video-MME [21] and MLVU [90] dataset. In summary, our paper makes the following contributions:
@@
(3) Long Video Assistant (LongVA): With long context transfer and UniRes, we developed LongVA that can perceive more than 200K visual tokens, achieving SoTA performance on the Video-MME and MLVU dataset.
@@ 2 Related Work
Recent years have witnessed significant progress in Video Question-Answering[3]. To accurately measure the progress of the video LMMs’ performance, researchers have developed various benchmarks encompassing a broad spectrum of tasks. These range from fundamental visual perception tasks such as activity recognition[84], concept detection [80], and counting [28], to more complex visual reasoning tasks including compositional [24], causal [76, 83, 81], and situated reasoning [74]. However, most of those benchmarks focus on short videos, lacking data and metrics to test LMMs’ capability over a long context. Inspired by the NIAH test [23] in the language model community, we proposed V-NIAH to benchmark LMMs’ ability over long visual inputs with the minimum overhead of data collection and human annotation. Several concurrent works also developed multimodal versions of the Needle-in-a-haystack test [73, 90, 63, 71]. However, they only measure on several hundreds of frames and lack a strong baseline to properly analyze the properties of visual context length.
@@ 3.2 Aligning Long Language Model Using Short Vision Data
Inspired by the AnyRes encoding scheme in LLaVA-NeXT [47, 35], we designed UniRes that provides a unified encoding scheme for both images and videos, as shown in Figure 2. Unlike AnyRes which retains a small base image and flattens ViT patches across the grids, UniRes removes the base image, flattens patches within each grid, and 2x2 pool the visual features by default (Appendix B). This approach allows us to maintain consistent representation when extending image data into videos where multiple frames are viewed as multiple grids in a row.
@@ 5 Experiments
We primarily assess the long visual capability of LongVA on two benchmarks: V-NIAH (Section 5.1 and Video-MME [21] (Section 5.2). V-NIAH provides quick signals about the visual context length of LongVA. However, it only tests the model’s ability to retrieve information and does not cover other abilities necessary for a real-world long video assistant. Therefore, we also include LongVA’s performance on Video-MME, a comprehensive evaluation suite for video LMMs that includes diverse data types and qualitative annotations. Video-MME is an ideal benchmark for assessing LMMs’ ability to handle long videos in real-world scenarios, given its average video duration of 1017 seconds and the inclusion of short, medium, and long subsets. We further include the benchmark results on MLVU[90] in Appendix C.
@@ 7 Conclusion
8 Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2022-01-029). Besides, this project is supported by NTU NAP, MOE AcRF Tier 2 (MOE-T2EP20221-0012), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as advise from the industry partner(s).
@@ Appendix A Needle In A Haystack Test
Appendix B UniRes Encoding Scheme
Appendix C MLVU Results
Methods Input Holistic Single Detail Multi Detail Averages
TR AR VS NQA ER PQA SSC AO AC M-Avg G-Avg
GPT-4o [55] 0.0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
LLaVA-1.6 [47] 0016 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
InternVL-1.5 [14] 0016 frm 78.8 67.0 3.16 52.7 43.5 54.4 4.88 32.8 23.8 50.4 4.02
MovieChat [64] 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
TimeChat [60] 0096 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
LLaMA-VID [41] 0001 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
MA-LMM [25] 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
ShareGPT4Video [12] 0016 frm 75.8 51.5 2.52 47.6 43.2 48.4 5.02 34.0 23.3 46.4 3.77
VideoChat2_HD [39] 0016 frm 77.3 60.5 3.38 46.2 48.9 50.1 4.59 23.2 29.1 47.9 3.99
VideoLlaMA2 [15] 0016 frm 74.6 64.5 2.79 49.9 43.8 45.1 5.18 34.0 27.4 48.5 3.99
LongVA (ours) 0256 frm 83.3 58.5 3.39 69.3 50.0 67.2 5.26 38.6 27.2 56.3 4.33
Table 8: Evaluation results by the authors of MLVU [90]. The highest scores excluding GPT-4o are bolden. TR: Topic Reasoning. AR: Anomaly Recognition. VS: Video Summary; NQA: Needle QA; ER: Ego Reasoning; PQA: Plot QA; SSC: Sub-Scene Captioning; AO: Action Order; AC: Action Count; M-Avg: the average performance of multiple-choice tasks; G-Avg: the average performance of generation tasks.
Table 8 includes the evaluation results by the authors of MLVU [90] on their benchmark. LongVA achieves state-of-the-art results among open-source models and is only second to GPT-4o.
Appendix D Visual Needle In A Haystack Test